OpenStudy (anonymous):

Express the complex number in trigonometric form. 3 - 3i

OpenStudy (ikram002p):

z=x+ iy \(z=r ( \cos \theta + i \sin \theta ) \) \(r =\sqrt {x^2+y^2}\) \(x= r \cos \theta \) \(y= r \cos \theta \) 

OpenStudy (ikram002p):

i made a typo \(y=r \sin \theta\)

OpenStudy (anonymous):

\[r=\sqrt{3^2+(-3)^2}=\sqrt{18}=3\sqrt2\] \[3=3\sqrt2\cos\theta~~\Rightarrow~~\cos\theta=\frac{1}{\sqrt2}~~\Rightarrow~~\theta=\frac{\pi}{4},\text{ or }\frac{7\pi}{4}\] \[-3=3\sqrt2\sin\theta~~\Rightarrow~~\sin\theta=-\frac{1}{\sqrt2}~~\Rightarrow~~\theta=\frac{5\pi}{4},\text{ or }\frac{7\pi}{4}\] Clearly the angle must be \(\dfrac{7\pi}{4}\), so \[3-3i=3\sqrt2\left(\cos\frac{7\pi}{4}+i\sin\frac{7\pi}{4}\right)\]